OpenStudy (anonymous):

On the question of =When is the Rational Expression Undefined, How do I solve these problems: 3/x ; 4/2-x; 5/xsquared-25; x-4/xsquared-4x+3; 5/x squared+1 Please give clear, concise directions and explanations Please do NOT send me youtube tutorials; I have BSOD (BlueScreenOfDeath), and playing videos causes it to happen. Thank you.

OpenStudy (anonymous):

You want to find when the bottom part, the denominator, is equal to 0. Since you can not divide by 0, the expression is undefined at those points.

OpenStudy (anonymous):

If the denominator is a linear equation, just solve for x. If it is a quadratic equation, use the quadratic formula to find when it is equal to 0.

OpenStudy (anonymous):

For example\[\frac{5}{x^2-25} \]Denominator is \(x^2-25\). So we find\[ x^2-25=0 \]In this case we don't even need quadratic formula, since we have a difference of squares and those are easy to factor \[ (x-5)(x+5)=0 \]So this equation is satisfied when \(x=5 \) or \(x=-5\). That means the expresson \(\large \frac{5}{x^2-25}\) is undefined at those values of \(x\).

OpenStudy (anonymous):

for the 5th problem I put up :\[\frac{ 5 }{ x ^{2}+1 }\] There is no way to factor for that.

OpenStudy (anonymous):

Which means that is is always defined.

OpenStudy (anonymous):

You'd need a complex number to get it equal to 0, but for these problems we tend to stick to just real numbers.

OpenStudy (anonymous):

Alright, on a different question, I was given a product(-5) and a sum(-4) I have to write the quadratic equation using the formula \[ax^2+bx+c\] How do I know which terms are a b and c, and how do I know where to place them